OpenStudy (anonymous):

Last one of the day :) sqr root sign with a 54 inside it I have to simplify by factoring I came up with 3sqr root 6 is this right? I wish I knew how you typed out all the symbols on here so I could do show the problem properly LOL @satellite73

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

Thanks! :) As always I appreciate all your time.

OpenStudy (anonymous):

\[54=9\times 6\] \[\sqrt{54}=\sqrt{9\times 6}=\sqrt{9}\sqrt{6}=3\sqrt6\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

do It put the 3 in front of the square root or as an exponent ..so i don't write it wrong.

OpenStudy (anonymous):

you put the three out front

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw again